An aluminum wire of diameter D and length L has a resistance of 1.5 Ω. Find the length of another aluminum wire at the same temperature if its diameter is 2D and its resistance is 6 Ω.


1
Expert's answer
2021-05-18T07:13:02-0400

Wire of the first case , Aluminum has the following information .

Length = L

Resistance = 1.5 "\Omega"

Diameter = D

Area A= "\pi r^2= \pi (\frac{D}{2})^2= \pi \frac{D^2}{4}"


R= "\rho \frac{L}{A}" , "\rho= \frac{AR}{L}"


"\rho" = "\frac{\pi \times 0.25D^2\times 1.5}{L} = \frac{0.375D^2}{L} \pi"

"\rho" = Resistivity is constant for all size of one type of material at a particular temperature .


Therefore the second aluminum also has the same resistivity as="\frac{0.375D^2}{L} \pi"


the second aluminum piece has the following information,


R= 6 "\Omega" Diameter = 2D ,

Area = "\pi r^2= \pi (\frac{2D}{2})^2 = \pi D^2"


Length of the wire = "\frac{AR}{\rho}" = "\pi(\frac{D^2 \times 6\times L}{\pi \times 0.375\times D^2})= 16 L"